OpenStudy (anonymous):

solve the equation for x: 7.63^(2x) = 9.18^(x+1) any idea how to?

OpenStudy (amistre64):

take the log of each side

OpenStudy (amistre64):

log(L^(2x)) = log(R^(x+1)) 2x log(L) = (x+1) log(R) 2x log(R) ----- = ----- x+1 log(L)

OpenStudy (amistre64):

2x ----- = 1.09 x+1 2x = 1.09 (x+1) 2x = 1.09x + 1.09 ... can you get it from here?

OpenStudy (anonymous):

yeah sure. Thank you so much :)